a rectangle has an area of 84 square cm. what are the possible dimensions?
Answer by Alan3354(69443)

a rectangle has an area of 84 square cm. what are the possible dimensions?
--------------
Area = L*W = 84
L = 84/W
-----
0 < W < 84
Pick any number for W and find L
W = 1
W = 2
W = 3
W = pi
W = 0.001
etc